# Commute Graph Neural Networks

### v4

@@ 5.2 Similarity-based Graph Rewiring
Both the fundamental matrix defined in Eq. 3 and Weighted DiLap requires 4.1 to ensure the existence and uniqueness of the Perron vector $\pi$, conditions that are not universally met in general graphs. To fulfill the irreducibility and aperiodicity assumptions,  (Tong et al., 2020a) introduce a teleporting probability uniformly distributed across all nodes. This method, inspired by PageRank (Page et al., 1999), amends the transition matrix to $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, where $\gamma\in(0,1)$. $\mathbf{P}_{pr}$ allows for the possibility that a random walker may choose a non-neighbor node for the next step with a probability of $\frac{1-\gamma}{N}$. This adjustment ensures that $\mathbf{P}_{pr}$ is irreducible and aperiodic, so it has a unique $\pi$. However, this approach leads to a complete graph represented by a dense matrix $\mathbf{P}_{pr}$, posing significant challenges for subsequent computational processes.
@@
Rather than employing $\mathbf{P}_{pr}$ as the transition matrix, we introduce a graph rewiring method based on feature similarity to make a given graph irreducible, while maintaining the sparsity. As outlined in Proposition 1, to transform the digraph into a strongly connected structure, it is essential that each node possesses a directed path to every other node. To this end, we initially construct a simple and irreducible graph $G^{\prime}$ with all $N$ nodes, then add all edges from $G^{\prime}$ the original digraph $G$, thereby ensuring $G$’s irreducibility. The construction of $G^{\prime}$ begins with the calculation of the mean of node features as the anchor vector $\bm{a}$. Then we determine the similarity between each node and the anchor, sort the similarity values, and return the sorted node indices, denoted as $S\in\mathbb{R}^{N}$:
$\bm{a}=\frac{\sum_{i}\mathbf{X}_{i}}{N},\quad s_{i}=\cos(\bm{a},\mathbf{X}_{i}),\quad S=\arg\mathrm{sort}(\{s_{i}\}_{i=1}^{N})$ (7)
@@
where $\cos(\bm{a},\mathbf{X}_{i})$ is the cosine similarity between node features of $v_{i}$ and $\bm{a}$, and $\mathrm{argsort}(\cdot)$ yields the indices of nodes that sort similarity values $\{s_{i}\}_{i=1}^{N}$. We then connect the nodes one by one with undirected (bidirectional) edges following the order in $S$ to construct $G^{\prime}$, as shown in Fig. 2. Given that $G^{\prime}$ is strongly connected, adding all its edges into $G$ results in a strongly connected digraph $\widetilde{G}$, which is irreducible. To achieve aperiodicity, self-loops are further added to $\widetilde{G}$.
@@
Figure 2: The sorted node indices in $S$ are connected one by one with undirected edges to construct $G^{\prime}$, then adding all edges from $G^{\prime}$ to $G$ to generate $\widetilde{G}$.
@@
This rewiring approach satisfies 4.1 and maintains graph sparsity. Additionally, adding edges between nodes with similar features only minimally alters the overall semantics of the original graph. Based on $\widetilde{G}$ and its corresponding $\widetilde{\mathbf{P}}$ and $\widetilde{\mathbf{\Pi}}$, we have the deterministic Weighted DiLap $\widetilde{\mathcal{T}}$.
@@ 5.3 From DiLap to Commute Time
Given a rewired graph $\widetilde{G}$, the Weighted DiLap is defined as $\widetilde{\mathcal{T}}=\widetilde{\mathbf{\Pi}}(\widetilde{\mathbf{D}}^{-1}-\widetilde{\mathbf{P}})$. Then the fundamental matrix $\mathbf{Z}$ of $\widetilde{G}$ can be solved by:
$\mathbf{Z}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\mathcal{R}^{\dagger}\widetilde{\mathbf{\Pi}}^{\frac{1}{2}},$ (8)
where $\mathcal{R}=\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}\widetilde{\mathcal{T}}\widetilde{\mathbf{\Pi}}^{-\frac{1}{2}}-\widetilde{\mathbf{D}}^{-1}+\mathbf{I}$ and the superscript $\dagger$ means Moore–Penrose pseudoinverse of the matrix.
The proof is given in Section A.2. Leveraging Lemma 5.1 and using Eq. 4, we can further compute the hitting times and commute times in terms of $\widetilde{\mathcal{T}}$ with the following theorem.
@@
$h(v_{i},v_{j})$ $=\frac{\mathcal{R}^{\dagger}_{jj}}{\pi_{j}}-\frac{\mathcal{R}^{\dagger}_{ij}}{\sqrt{\pi_{i}\pi_{j}}},$ (9)
$c(v_{i},v_{j})$ $=h(v_{i},v_{j})+h(v_{j},v_{i})=\frac{\mathcal{R}^{\dagger}_{jj}}{\pi_{j}}+\frac{\mathcal{R}^{\dagger}_{ii}}{\pi_{i}}-\frac{\mathcal{R}^{\dagger}_{ij}}{\sqrt{\pi_{i}\pi_{j}}}-\frac{\mathcal{R}^{\dagger}_{ji}}{\sqrt{\pi_{i}\pi_{j}}}.$
@@
$\mathcal{H}=(e\otimes\pi^{-1})(\mathcal{R}^{\dagger}\odot\mathbf{I})-\mathcal{R}^{\dagger}\odot(\pi^{-\frac{1}{2}}\otimes\pi^{-\frac{1}{2}}),\quad\mathcal{C}=\mathcal{H}+\mathcal{H}^{\top}$ (10)
where $\odot$ denotes Hadamard product, and $\otimes$ is outer product. $\mathcal{H}_{ij}$ and $\mathcal{C}_{ij}$ correspond to the hitting and commute time from $v_{i}$ to $v_{j}$ respectively. The computation of commute times via DiLap, in contrast to the method delineated in Theorem 4.2, is primarily motivated by efficiency concerns. Specifically, Eq. 3 necessitates the inversion of a dense matrix with complexity $\mathcal{O}(N^{3})$, whereas our DiLap-based method hinges on computing the pseudoinverse of a sparse matrix $\mathcal{R}$. The pseudoinverse of $\mathcal{R}$ can be efficiently determined using SVD. Given the sparse nature of $\mathcal{R}$, we can employ well-established techniques such as the randomized truncated SVD algorithm (Halko et al., 2011; Cai et al., 2023), which takes advantage of sparsity, to reduce the time complexity to $\mathcal{O}(q|E|)$, where $|E|$ denotes the number of edges reflecting the sparsity (See Section A.3). Next, we present Commute Graph Neural Networks (CGNN) based on $\mathcal{C}$.

### v5

@@ 4.2 Similarity-based Graph Rewiring
Both the fundamental matrix defined in Eq. 3 and Weighted DiLap requires 3.1 to ensure the existence and uniqueness of the Perron vector $\pi$, conditions that are not universally met in general graphs. To fulfill the irreducibility and aperiodicity assumptions,  Tong et al. (2020a) introduce a teleporting probability uniformly distributed across all nodes. This method, inspired by PageRank (Page et al., 1999), amends the transition matrix to $\mathbf{P}_{pr}=\gamma\mathbf{P}+(1-\gamma)\frac{ee^{\top}}{N}$, where $\gamma\in(0,1)$. $\mathbf{P}_{pr}$ allows for the possibility that a random walker may choose a non-neighbor node for the next step with a probability of $\frac{1-\gamma}{N}$. This adjustment ensures that $\mathbf{P}_{pr}$ is irreducible and aperiodic, so it has a unique $\pi$. However, this approach leads to a complete graph represented by a dense matrix $\mathbf{P}_{pr}$, posing significant challenges for subsequent computational processes.
@@
Rather than employing $\mathbf{P}_{pr}$ as the transition matrix, we introduce a graph rewiring method based on feature similarity to make a given graph irreducible, while maintaining the sparsity. As outlined in Proposition 3.2, to transform the digraph into a strongly connected structure, it is essential that each node possesses a directed path to every other node. To this end, we initially construct a simple and irreducible graph $G^{\prime}$ with all $N$ nodes, then add all edges from $G^{\prime}$ to the original digraph $G$, thereby ensuring $G$’s irreducibility. The construction of $G^{\prime}$ begins with the calculation of the mean of node features as the anchor vector $\boldsymbol{a}$. Then we determine the similarity between each node and the anchor, sort the similarity values, and return the sorted node indices, denoted as $\Omega\in\mathbb{R}^{N}$:
$\boldsymbol{a}=\frac{\sum_{i}\mathbf{X}_{i}}{N},\quad\omega_{i}=\cos(\boldsymbol{a},\mathbf{X}_{i}),\quad S=\arg\mathrm{sort}(\{\omega_{i}\}_{i=1}^{N})$ (7)
@@
where $\cos(\boldsymbol{a},\mathbf{X}_{i})$ is the cosine similarity between node features of $v_{i}$ and $\boldsymbol{a}$, and $\mathrm{argsort}(\cdot)$ yields the indices of nodes that sort similarity values $\{\omega_{i}\}_{i=1}^{N}$. We then connect the nodes one by one with undirected (bidirectional) edges following the order in $S$ to construct $G^{\prime}$, as shown in Figure 2. Given that $G^{\prime}$ is strongly connected, adding all its edges into $G$ results in a strongly connected digraph $\widetilde{G}$, which is irreducible. To achieve aperiodicity, self-loops are further added to $\widetilde{G}$.
@@
Figure 2: The sorted node indices in $\Omega$ are connected one by one with undirected edges to construct $G^{\prime}$, then adding all edges from $G^{\prime}$ to $G$ to generate $\widetilde{G}$.
@@
This rewiring approach satisfies 3.1 and maintains graph sparsity. Additionally, adding edges between nodes with similar features only minimally alters the overall semantics of the original graph. Based on $\widetilde{G}$ and its corresponding $\widetilde{\mathbf{P}}$, $\widetilde{\mathbf{B}}$, and $\widetilde{\mathbf{\Pi}}$, we have the deterministic Weighted DiLap $\widetilde{\mathcal{T}}$.
@@ 4.3 From DiLap to Commute Time
Given a rewired graph $\widetilde{G}$, the Weighted DiLap is defined as $\widetilde{\mathcal{T}}=\widetilde{\mathbf{\Pi}}\widetilde{\mathbf{B}}\mathrm{diag}\left(\left\{\widetilde{\mathbf{P}}_{ij}\right\}_{(v_{i},v_{j})\in E}^{M}\right)\widetilde{\mathbf{B}}^{\top}$. Then the fundamental matrix $\mathbf{Z}$ of $\widetilde{G}$ can be solved by:
$\mathbf{Z}=\widetilde{\mathcal{T}}^{\dagger}\widetilde{\mathbf{\Pi}}=\widetilde{\mathbf{T}}^{\dagger},$ (8)
where the superscript means Moore–Penrose pseudoinverse of the matrix.
Leveraging Lemma 4.2 and using Eq. 4, we can further compute the hitting times and commute times in terms of $\widetilde{\mathcal{T}}$ with the following theorem.
@@
$h(v_{i},v_{j})$ $=\frac{\widetilde{\mathbf{T}}^{\dagger}_{jj}}{\pi_{j}}-\frac{\widetilde{\mathbf{T}}^{\dagger}_{ij}}{\sqrt{\pi_{i}\pi_{j}}},$ (9)
$c(v_{i},v_{j})$ $=\frac{\widetilde{\mathbf{T}}^{\dagger}_{jj}}{\pi_{j}}+\frac{\widetilde{\mathbf{T}}^{\dagger}_{ii}}{\pi_{i}}-\frac{\widetilde{\mathbf{T}}^{\dagger}_{ij}}{\sqrt{\pi_{i}\pi_{j}}}-\frac{\widetilde{\mathbf{T}}^{\dagger}_{ji}}{\sqrt{\pi_{i}\pi_{j}}}.$
@@
$\mathcal{H}$ $=(e\otimes\pi^{-1})(\widetilde{\mathbf{T}}^{\dagger}\odot\mathbf{I})-\widetilde{\mathbf{T}}^{\dagger}\odot(\pi^{-\frac{1}{2}}\otimes\pi^{-\frac{1}{2}})$ (10)
$\mathcal{C}$ $=\mathcal{H}+\mathcal{H}^{\top}$
where $\odot$ denotes Hadamard product, and $\otimes$ is outer product. $\mathcal{H}_{ij}$ and $\mathcal{C}_{ij}$ correspond to the hitting and commute time from $v_{i}$ to $v_{j}$ respectively. The computation of commute times via DiLap, in contrast to the method delineated in Theorem 3.3, is primarily motivated by efficiency concerns. Specifically, Eq. 3 necessitates the inversion of a dense matrix with complexity $\mathcal{O}(N^{3})$, whereas our DiLap-based method hinges on computing the pseudoinverse of a sparse matrix $\widetilde{\mathbf{T}}$. The pseudoinverse of $\widetilde{\mathbf{T}}$ can be efficiently determined using SVD. Given the sparse nature of $\widetilde{\mathbf{T}}$, we can employ well-established techniques such as the randomized truncated SVD algorithm (Halko et al., 2011; Cai et al., 2023), which takes advantage of sparsity, to reduce the time complexity to $\mathcal{O}(q|E|)$, where $|E|$ denotes the number of edges reflecting the sparsity (See Section A.4). Next, we present CGNN based on $\mathcal{C}$.